# Piscataway, NJ 08854 USA

### v1

@@ 1.4 Tensor fields
Two aspects of the proposal in hilbertbundles ; hilbertbundles1 have not been derived from AdS/CFT models. The first is the claim that the near horizon CFT is constructed from free fermions in one to one correspondence with the (cut-off) spectrum of eigenspinors of the Dirac operator on the holographic screen of the diamond. The second is that the CFT should actually be an interacting theory, in order to capture the fast scrambling property of horizons. The absence of the second property is clearly related to our study of the strict $N=\infty$ limit. The boundary CFT is an integrable system in that limit, and cannot exhibit fast scrambling since all connected $k>2$ point functions vanish. Fast scrambling should appear once $1/N$ corrections are taken into account properly. At any rate, it now appears that, within the context of the Carlip-Solodukhin ansatz, fast scrambling is implemented by marginally irrelevant perturbations of the free fermion CFT, rather than an interacting CFT toappear .
@@ Acknowledgements.
The work of S. A is supported in part by the DOE under grant DE-SC0010008.

### v2

@@ 1.4 Tensor fields
Three aspects of the proposal in hilbertbundles ; hilbertbundles1 have not been derived from AdS/CFT models. The first is the claim that the near horizon CFT is constructed from free fermions in one to one correspondence with the (cut-off) spectrum of eigenspinors of the Dirac operator on the holographic screen of the diamond. The second is that the CFT should actually be an interacting theory, in order to capture the fast scrambling property of horizons. The third, and the most important, is that the CFT obtained through tensor network approximation with the large N formalism of AdS/CFT fails to be a CFT with a gravitational dual.
The absence of the second property is clearly related to our study of the strict $N=\infty$ limit. The boundary CFT is an integrable system in that limit, and cannot exhibit fast scrambling since all connected $k>2$ point functions vanish. Fast scrambling should appear once $1/N$ corrections are taken into account properly. At any rate, it now appears that, within the context of the Carlip-Solodukhin ansatz, fast scrambling is implemented by marginally irrelevant perturbations of the free fermion CFT, rather than an interacting CFT toappear .
@@
However, the most serious issue with our “derivation” of the Carlip-Solodukhin prescription from the standard large N formalism of AdS/CFT combined with tensor networks, is that we are not able to account for the large value of the central charge cc [^4] This was pointed out to one of us (T.B.) by D. Harlow when the material in this paper was presented at an informal seminar at MIT.. The reason for this is that for holographic CFTs, models with an Einstein-Hilbert dual, there are low energy states of the emergent Minkowski scattering theory in the central node of the network, which do not correspond to limits of low energy states in the CFT. This was pointed out in tbwfads2 , where it was argued that the indefinitely large number of soft gravitons in flat space scattering amplitudes arose from multipoint connected correlators in the CFT, in which most of the boundary operators create bulk excitations that do not enter the Polchinski-Susskind “arena” diamond, whose boundary converges to conformal infinity in Minkowski space. The surviving states, which appear to a detector on the central geodesic as low energy states living on the boundary of the arena, are, from the boundary CFT point of view, complicated linear combinations of energy eigenstates, including high energy states. Recall that high energy scattering can produce large number of low energy gravitons.
Thus, the simple lesson extracted from the work of Evenbly and Vidal from soluble CFTs, that the spectrum of small lattices in a tensor network approximation to a CFT is a good approximation to the low lying spectrum of the CFT, does not really extend to CFTs with Einstein-Hilbert duals. The bulk of low energy states on small lattices are soft graviton states, which arise from complicated entanglement patterns with multipoint correlators outside the central node. Furthermore the entropy of the central node comes from terms that are essentially non-perturbative in the $1/N$ expansion, so it does not make sense to think of it as being captured by the $1/N$ corrections.
@@ Acknowledgements.
The work of S. A is supported in part by the DOE under grant DE-SC0010008. T.B. thanks D. Harlow for an illuminating comment on this material during a talk at MIT.